# ReDeLEx: A Framework for Relational Deep Learning Exploration

### v1

@@ 4.3.1 Related Work
As outlined in the Introduction, the ReDeLEx framework builds upon the CTU relational dataset collection [29] which it integrates with the RelBench [33] interface to facilitate a wider scope of RDL [15] experimentation. As such, it is naturally related to recent works introducing new RDL models, which include [32, 43, 7]. Besides RDL, related work includes other dataset and benchmarking frameworks that address some facets of learning from relational data, including [39, 40]. The most salient feature of ReDeLEx, within the context of related work, is the provided bridge between the traditional relational learning methods [14] and the contemporary RDL [15].
@@ Appendix 0.A Experimental Setup
All deep learning models (including RDL) were trained for a minimum of 10 epochs with a total minimum of 1000 steps of Adam [23] optimizer with a learning rate of 0.001. Experiments on the RDL models were conducted with fixed hyperparameters with two exceptions—the number of layers of the graph neural model and the neighborhood graph sampling rate, which were searched for in a grid hyperparameter optimization. The hyperparameters include batch size—set to 512, message-passing aggregation function—set to summation, embedding vectors dimension, which is the same for both row and attribute embedding vectors—set to 64, neighborhood sampling rate—iterated over the values of 16, 32 and 64, and number of message-passing layers—a value in range of 1 to 4.

### v2

@@ 4.3.1 Related Work
As outlined in the Introduction, the ReDeLEx framework builds upon the CTU relational dataset collection [motl2015ctu] which it integrates with the RelBench [robinson2024relbench] interface to facilitate a wider scope of RDL [fey2024position] experimentation. As such, it is naturally related to recent works introducing new RDL models, which include [peleska_transformers_2024, zahradnik2023deep, chen_relgnn_2025]. Besides RDL, related work includes other dataset and benchmarking frameworks that address some facets of learning from relational data, including [vogel2024wikidbs, wang_4dbinfer_2024]. The most salient feature of ReDeLEx, within the context of related work, is the provided bridge between the traditional relational learning methods [Raedt] and the contemporary RDL [fey2024position].
@@ Appendix 0.A Experimental Setup
All deep learning models (including RDL) were trained for a minimum of 10 epochs with a total minimum of 1000 steps of Adam [Kingma2014] optimizer with a learning rate of 0.001. Experiments on the RDL models were conducted with fixed hyperparameters with two exceptions—the number of layers of the graph neural model and the neighborhood graph sampling rate, which were searched for in a grid hyperparameter optimization. The hyperparameters include batch size—set to 512, message-passing aggregation function—set to summation, embedding vectors dimension, which is the same for both row and attribute embedding vectors—set to 64, neighborhood sampling rate—iterated over the values of 16, 32 and 64, and number of message-passing layers—a value in range of 1 to 4.